OpenStudy (moonlitfate):

f(x) = 4-x^2 determine the lower sum s(n) over the interval [1,2]

OpenStudy (anonymous):

is this for a calculus class?

OpenStudy (moonlitfate):

Yes.

OpenStudy (anonymous):

ok, is this in a section of integration?

OpenStudy (moonlitfate):

Yes; the whole area under a curve kind of thing.

OpenStudy (anonymous):

well here is a graph of the function

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y+%3D+4-x%5E2+

OpenStudy (anonymous):

you want to integrate over the interval from 1 to 2.

OpenStudy (moonlitfate):

I all ready know what \[\Delta x = (2-1)/n = \frac{ 1 }{ n }\] And that the end points are going to be \[a + (i - 1)\Delta x = 1+ (\frac{ i-1 }{ n })\]

OpenStudy (moonlitfate):

We aren't supposed to find the limit. :/

OpenStudy (moonlitfate):

The sum has to be in terms of n, if that makes sense.

OpenStudy (anonymous):

my, bad I haven't taken calculus in a while, and after you get through the calc sequence you never use limits or find this stuff. you mostly evaluate integrals. however, this link may help you http://www.tc3.edu/instruct/sbrown/calc/ulsum.htm

OpenStudy (moonlitfate):

It's finding the sum using sigma notation. \[\sum_{i = 1}^{n}f(x _{i})\Delta x\]